OpenStudy (anonymous):

give the domain of the function F(x)=2/x^2+3

OpenStudy (anonymous):

Is it \[F(x)=2/(x^{2}+3)\]

OpenStudy (anonymous):

Then domain will be \[(-\infty,+\infty)\]

OpenStudy (anonymous):

its all real numbers?

OpenStudy (anonymous):

yes All real numbers

OpenStudy (anonymous):

how though? there are many more that i have to do

OpenStudy (anonymous):

am i trying to make it equal zero?

OpenStudy (anonymous):

Domain is the set or region where your F(x) is properly defined. now x^2+3 in the denominator(denominator should not be zero). For all real number it cannot be a zero. So, for all real numbers you will have F(x) properly defined x^2+3 will be zero for x=\[\sqrt{-3}\]. This is imaginary

OpenStudy (anonymous):

oh.. lol okay thank you